A 4.5V flashlight draws current from of 400mA when the light is on. If the light is on for 50 seconds, what quantity of electrical potential energy is converted into heat, using Given, Required, Analysis & Solution, Paraphrase.
1
Expert's answer
2012-06-12T09:53:55-0400
Q = dt · U · I = 50 · 4.5 · 0.4 = 90 J